# Anomalous nuclear spin coherence in superconducting Nb3Sn

### v1

@@ VI Acknowledgement
We thank Jim Sauls for helpful discussions and Michael Brown for providing Sample 1. We also acknowledge assistance from John Scott and Daehan Park in the ultra-low temperature laboratory. This work was supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract No. DE-AC02-07CH11359. A portion of this work was performed at the National High Magnetic Field Laboratory, which is supported by National Science Foundation Cooperative Agreement No. DMR-2128556 and the State of Florida. The research work at Florida State University was primarily funded by the US. Department of Energy, Office of High Energy Physics, DE-SC0012083.

### v3

@@ VI Acknowledgement
We thank Jim Sauls and Andy Mounce for helpful discussions and Michael Brown for providing Sample 1. We also acknowledge assistance from John Scott and Daehan Park in the ultra-low temperature laboratory. This work was supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract No. DE-AC02-07CH11359. A portion of this work was performed at the National High Magnetic Field Laboratory, which is supported by National Science Foundation Cooperative Agreement No. DMR-2128556 and the State of Florida. The research work at Florida State University was primarily funded by the US. Department of Energy, Office of High Energy Physics, DE-SC0012083.
VII SUPPLEMENTARY MATERIALS
VIII NMR Spectrum
Figure 5: NMR 93Nb spectrum of Sample 1. The field sweep spectrum showing raw data in blue and fitted to the theory for random oriented powder in red; seven components of the nine are evident where two of the satellites at highest and lowest frequencies are unresolved. The central component is evident with a resolved splitting as expected in second order perturbation theory in the electric field gradient. Similar spectra were measured for Sample 2.
IX Critical Field
Figure 6: Temperature dependence of the upper critical field of Nb3Sn. The upper critical field are evaluated with a 90%Rn criterion from transport measurements. The blue curve (Sample 1) and red curve (Sample 2) are calculated using Werthamer, Helfand and Hohenberg (WHH) fitting.
X SEM Image
Figure 7: SEM image of Sample 2, prior to ball milling. Note the uniform shade of what appear to be single crystal grains.
XI Resistance Measurement
Figure 8: Temperature dependence of the resistance at different magnetic fields for Sample 2.